OpenStudy (anonymous):

Help again yo!? Chuck is making a map for Geography. In order to draw the map, he must create a scale converting the measured inches on the map to actual miles. Use the table above to identify the scale he used for his map. A.1 in = 12 miles B.1 in = 6 miles C.1 in = 8 miles D. 1 in = 24 miles

OpenStudy (anonymous):

GRAPH Length in Inches 1 1/4 1 1/2 1 3/4 2 1/4 2 1/2 2 3/4 Length in Miles 30 36 42 54 60 66

OpenStudy (mathmath333):

\(\text{difference in inches=difference in miles}\\ 1\cfrac{1}{2}-1\cfrac{1}{4}~~inches=36-30~~miles\\ 1+\dfrac{1}{2}-1-\dfrac{1}{4}~~inches=6~~miles\\ \dfrac{1}{2}~~inches=6~~miles\)

OpenStudy (mathmath333):

*error \(\dfrac{1}{4}inches=6~~milles\)